# Purified pseudomode model for nonlinear system-bath interactions

### v1

@@ IV.2 Fluorescence spectrum of quantum dots in the presence of phonon environments
We next consider a more complex physical example, namely, the fluorescence spectrum of a two-level quantum dot in the presence of a phonon environment, which affects the excitation energy of the quantum dot, leading to the bath coupling operator $s=\sigma_{+}\sigma_{-}$. To measure the fluorescence spectrum, the quantum dot is driven by a resonant laser so that in the frame rotating with the laser frequency the bare Hamiltonian of the quantum dot is $H_{S}=\Omega\sigma_{x}/2$ where $\Omega$ is the Rabi frequency. The quantum dot also exhibits a radiative decay with rate $\kappa$, corresponding to a Lindblad dissipator with the collapse operator $\sigma_{-}$. We employ the super-Ohmic spectral density $J(\omega)=\alpha_{p}\omega^{3}\exp(-\omega^{2}/2\omega_{b}^{2})$ for the phonon environment, where $\alpha_{p}$ measures the coupling strength and $\omega_{b}$ is the cutoff frequency. The spectrum is obtained from the Fourier transform of the two-time correlation function, i.e.,
[…]
@@
[…]
@@ V Conclusions and outlook
As an outlook, the results presented in this paper could be applied for the modeling of other types of nonlinear system-bath interactions, such as the ones in Refs. [86, 87] which can be engineered in waveguide QED. Another possible future direction is to further extend the purified pseudomode model here to allow for the computation of bath properties [79].
@@ Acknowledgements.
We would like to thank Dr. W. Yang for interesting discussions about the role of nonlinearities in the pseudomode model. M.C. acknowledges support from NSFC (Grant No. 11935012) and NSAF (Grant No. U2330401). N.L. is supported by the RIKEN Incentive Research Program and by MEXT KAKENHI Grants No. JP24H00816 and No. JP24H00820.
@@ Appendix A Detailed derivations for the purification of the pseudomode model
$[\mathfrak{S},\mathfrak{L}_{0}]=-\mathfrak{V}+\mathfrak{V}^{\prime}+\mathfrak{V}_{\text{res}}+\mathcal{O}(a^{-1}),$ (41)
with $\mathfrak{L}_{0}=\mathfrak{L}_{S}+\mathfrak{L}_{\text{pm},0}^{a}+\big{(}\bar{s}\bar{X}_{+}^{n}-\tilde{s}^{\intercal}\tilde{X}_{-}^{n}\big{)}+\mathfrak{R}$. Here, the operator $\mathfrak{V}^{\prime}$ represents the emerging mediated interactions, and for consistency we require, not to contain any operators related to the modes $\tilde{d}_{l+}$, $\bar{d}_{l-}$, $\tilde{d}_{0}$ and $\bar{d}_{0}$, and to be $\mathcal{O}(a^{0})$. The operator $\mathfrak{V}_{\text{res}}$ contains terms ending with either $\tilde{X}_{+,d}$ or $\bar{X}_{-,d}$ or $\tilde{X}_{0,d}$ or $\bar{X}_{0,d}$, which, for the same reason concerning the term $\mathfrak{R}$ in Eq.(37) renders $\mathfrak{V}_{\text{res}}$ irrelevant to the dynamics. Importantly, the existence of the commutator in Eq. (41) ensures nested commutators appearing in the SW transformation (40), such as $[\mathfrak{S},\cdots,[\mathfrak{S},\mathfrak{L}_{0}]]$ with at least two $\mathfrak{S}$, or $[\mathfrak{S},\cdots,[\mathfrak{S},\mathfrak{V}]]$ with at least one $\mathfrak{S}$, to be at least of order $\mathcal{O}(a^{-1})$. Hence, in the limit $a\to+\infty$, they vanish in the expansion of the SW transformation (40), leading to the simple generator
@@
$\begin{array}[]{lll}\mathfrak{V}^{\prime}&=&\displaystyle\bar{s}\bigg{[}\tilde{X}_{-,d}^{n}+\sum_{r=1}^{n-1}\binom{n}{r}\bar{X}_{+}^{n-r}\tilde{X}_{-,d}^{r}\bigg{]}\\ &&\displaystyle-\tilde{s}^{\intercal}\bigg{[}\bar{X}_{+,d}^{n}+\sum_{r=1}^{n-1}\binom{n}{r}\bar{X}_{+,d}^{r}\tilde{X}_{-}^{n-r}\bigg{]},\\ \end{array}$ (46)
@@
$\begin{array}[]{lll}&\displaystyle id\rho_{\mathbf{m},\mathbf{n}}/dt\\ &\displaystyle=[H_{\text{S}},\rho_{\mathbf{m},\mathbf{n}}]\\ &\displaystyle+\sum_{l}\Big{(}(\nu_{l}-i\gamma_{l})m_{l}\rho_{\mathbf{m},\mathbf{n}}-(\nu_{l}+i\gamma_{l})n_{l}\rho_{\mathbf{m},\mathbf{n}}\Big{)}\\ &\displaystyle+\sum_{\mathbf{m}^{\prime}}\langle\mathbf{m}|X_{+}^{n}|\mathbf{m}^{\prime}\rangle s\rho_{\mathbf{m}^{\prime},\mathbf{n}}-\sum_{\mathbf{n}^{\prime}}\langle\mathbf{n}^{\prime}|X_{-}^{n}|\mathbf{n}\rangle\rho_{\mathbf{m},\mathbf{n}^{\prime}}s\\ &\displaystyle+\sum_{\mathbf{n}^{\prime}}\langle\mathbf{n}^{\prime}|X_{-,d^{\dagger}}^{n}|\mathbf{n}\rangle s\rho_{\mathbf{m},\mathbf{n}^{\prime}}-\sum_{\mathbf{m}^{\prime}}\langle\mathbf{m}|X_{+,d}^{n}|\mathbf{m}^{\prime}\rangle\rho_{\mathbf{m}^{\prime},\mathbf{n}}s\\ &\displaystyle+\sum_{r=1}^{n-1}\binom{n}{r}\sum_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}\bigg{[}\langle\mathbf{m}|X_{+}^{n-r}|\mathbf{m}^{\prime}\rangle\langle\mathbf{n}^{\prime}|X_{-,d^{\dagger}}^{n-r}|\mathbf{n}\rangle s\rho_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}\\ &\displaystyle-\langle\mathbf{m}|X_{+,d}^{n-r}|\mathbf{m}^{\prime}\rangle\langle\mathbf{n}^{\prime}|X_{-}^{n-r}|\mathbf{n}\rangle\rho_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}s\bigg{]},\end{array}$ (48)

### v2

@@ IV.2 Fluorescence spectrum of quantum dots in the presence of phonon environments
We next consider a more complex physical example, namely, the fluorescence spectrum of a two-level quantum dot in the presence of a phonon environment, which affects the excitation energy of the quantum dot, leading to the bath coupling operator $s=\sigma_{+}\sigma_{-}$. Unlike the previous example, the phonon environment consists of a continuum of bath modes, rendering exact results inaccessible. To measure the fluorescence spectrum, the quantum dot is driven by a resonant laser so that in the frame rotating with the laser frequency the bare Hamiltonian of the quantum dot is $H_{S}=\Omega\sigma_{x}/2$ where $\Omega$ is the Rabi freqeunce, chosen as $1\text{\,}\mathrm{meV}$ here to be consistent with Ref. [102]. The quantum dot also exhibits a radiative decay with rate $\kappa$, which can be accounted for by a Lindblad dissipator with the collapse operator $\sigma_{-}$, i.e., $\kappa(2\sigma_{-}\rho_{d}\sigma_{+}-\sigma_{+}\sigma_{-}\rho_{d}-\rho_{d}\sigma_{+}\sigma_{-})$ where $\rho_{d}$ denotes the dot state. Photons emitted via this radiative channel are then collected to measure the dot fluorescence spectrum. The last ingredient of this model is the phonon environment, here characterized by the super-Ohmic spectral density $J(\omega)=\alpha_{p}\omega^{3}\exp(-\omega^{2}/2\omega_{b}^{2})$, where $\alpha_{p}$ measures the coupling strength and $\omega_{b}$ is the cutoff frequency. The dot fluorescence spectrum is obtained from the Fourier transform of the two-time correlation function, i.e.,
$S_{\text{rf}}(\omega)=\text{Re}\,\int_{0}^{\infty}d\tau e^{i\omega\tau}\big[\langle\sigma^{+}(\tau)\sigma^{-}\rangle_{ss}-\langle\sigma^{+}(\infty)\sigma^{-}\rangle_{ss}\big],$ (28)
@@
In Fig. 3(a), we plot the power spectrum $S(\omega)$ of the phonon environment, related to the spectral density $J(\omega)$ via
$S(\omega)=(1+\coth(\beta\omega/2))\left[J(\omega)\theta(\omega)-J(-\omega)\theta(-\omega)\right],$ (29)
where $\theta(\omega)$ is the Heaviside step function. We have chosen the parameters $\omega_{b}=$1\text{\,}\mathrm{meV}$\$, $\alpha_{p}/(2\pi)^{2}=$0.002\text{\,}\mathrm{p}\mathrm{s}^{2}$\$ and $T=$4\text{\,}\mathrm{K}$\$ consistent with those in Ref. [102]. The spectral decompositions in Eqs. (15) and (16) are generated with the AAA algorithm [103, 82], which aims to find the rational approximation of a general function defined on the real axis, such as $S(\omega)$, in the barycentric representation
$S_{\text{AAA}}(\omega)=\sum_{l=1}^{M}\frac{z_{l}S(\omega_{l})}{\omega-\omega_{l}}\Big/\sum_{l=1}^{N}\frac{z_{l}}{\omega-\omega_{l}}.$ (30)
Here, $M\geq 1$ is an integer giving the order of the rational appproximation, $z_{l}\in\mathbb{C}$ are free parameters, and $\{\omega_{l}\in\mathbb{R}|l=1,\cdots,M\}$ is a set of support points, over which we have $S(\omega_{l})=S_{\text{AAA}}(\omega_{l})$ using the residue theorem. After the barycentric representation $S_{\text{AAA}}(\omega)$ is obtained, its poles $p_{l}$ in the lower half complex plane and the corresponding residues $r_{l}$ can be solved to yield the spectral decomposition in Eq. (15) with the parameters given by $w_{l}=-ir_{l}$, $\nu_{l}=\text{Re}\,p_{l}$ and $\gamma_{l}=-\text{Im}\,p_{l}$. Note that, in practice one often needs only a small portion NN ($N\leq M$) of these poles to achieve high approximation precision. The approximate power spectrum of the phonon bath is plotted in dashed line in Fig. 3(a). For the parameters chosen here, $N=4$ exponentials already produce a faithful approximation. In the two insets, we plot the fitting error $\epsilon_{S}=\int d\omega\lvert S(\omega)-S_{\text{AAA}}(\omega)\rvert$ versus $N$ (upper inset) and the fitting error $\epsilon_{C}=\lvert C(t)-C_{\text{AAA}}(t)\rvert$ as a function of time $t$ for $N=4$ (lower inset), where $C_{\text{AAA}}(t)$ is obtained from the Fourier transform of $S_{\text{AAA}}(\omega)$. As expected, $\epsilon_{S}$ decreases quickly when $N$ is increased while the main fitting error of $C(t)$ concentrates around $\omega=0$. A step-by-step recipe for constructing the purified pseudomode model for this example is summarized in detail in Appendix. C.
In Fig. 3(b), we show the dot fluorescence spectra in the absence of the phonon environment (blue line) and in the presence of the phonon environment with a linear (orange) or linear+quadratic (green) coupling. In the first case, the dot fluorescence spectrum displays the characteristic Mollow triplet [104] with two symmetric sidebands located around $\omega=\pm\Omega$. The emergence of sidebands can be explained in terms of laser-dressed states [104]. A linearly coupled phonon environment makes the dot fluorescence spectrum asymmetric and slightly broadens the three peaks, as shown by the orange line in Fig. 3(b). Asymmetric sidebands observed here arise from phonon-assisted transitions [105, 106, 107] between laser-dressed states and can be further attributed to the asymmetry of $S(\omega)$ between phonon emission (positive frequencies) and absorption (negative frequencies) processes, see Fig. 3(a). In the presence of a small quadratic correction, we observe an overall lift of the dot fluorescence spectrum compared to the pure linear case, see the green line in Fig. 3(b).
@@ V Conclusions and outlook
Let us summarize the improvement provided by our method with respect to previous studies. First, it offers a general recipe for modeling nonlinear system-bath interactions in the form of Eq. (2). This capability comes from the natural one-to-one correspondence of multi-time correlation functions between the original bath and the corresponding effective one made of pseudomodes, which avoids the combinatorial complexity in re-summing the Wick contractions required by other methods. For comparison, Refs. [80, 81] analyzed a similar problem within the HEOM formalism, but restricted to the case of linear interaction plus quadratic corrections, where the Wick re-summation is still tractable. Second, our method is non-perturbative and accounts for the multi-mode nature of bosonic quantum environments. In contrast, Ref. [108] simplified the description of the phonon bath by replacing it with a single, most relevant, bath mode, while Refs. [86, 93, 88] made use of the rotating wave approximation valid only for weak interaction strength in the context of cavity or waveguide QED. Our method thus paves the way for an accurate description of the effects of nonlinear light- and phonon-matter interactions. Finally, purification of the pseudomode state leads to a reduction in the local dimension of the state space, which can be exploited in tensor-network simulations to reduce computational complexity. This has been demonstrated in our previous work [79] for linear system-bath interactions.
As an outlook, the method developed in this work could be applied to the modeling of other types of nonlinear system-bath interactions, such as the one in Ref. [88] engineered in waveguide QED and the Peierls-type light-matter interactions realized in cavity quantum materials [109, 110, 111]. Possible further extensions of this method include the modeling of nonlinear interactions in fermionic environments [112], the computation of bath properties [79, 113], and the incorporation of stochastic techniques [73] and tensor-network algorithms [114, 115, 116].
@@ Acknowledgements.
We would like to thank Dr. W. Yang for interesting discussions about the role of nonlinearities in the pseudomode model. M.C. acknowledges support from NSFC (Grant No. 12088101) and NSAF (Grant No. U2330401).
@@ Appendix A Detailed derivations for the purification of the pseudomode model
$\mathfrak{S},\mathfrak{L}_{0}]=-\mathfrak{V}+\mathfrak{V}^{\prime}+\mathfrak{V}_{\text{res}}+\mathcal{O}(a^{-1}),$ (45)
with $\mathfrak{L}_{0}=\mathfrak{L}_{S}+\mathfrak{L}_{\text{pm},0}^{a}+\big(\bar{s}\bar{X}_{+}^{n}-\tilde{s}^{\intercal}\tilde{X}_{-}^{n}\big)+\mathfrak{R}$. Here, the operator $\mathfrak{V}^{\prime}$ represents the emerging mediated interactions, and for consistency we require, not to contain any operators related to the modes $\tilde{d}_{l+}$, $\bar{d}_{l-}$, $\tilde{d}_{0}$ and $\bar{d}_{0}$, and to be $\mathcal{O}(a^{0})$. The operator $\mathfrak{V}_{\text{res}}$ contains terms ending with either $\tilde{X}_{+,d}$ or $\bar{X}_{-,d}$ or $\tilde{X}_{0,d}$ or $\bar{X}_{0,d}$, which, for the same reason concerning the term $\mathfrak{R}$ in Eq.(41) renders $\mathfrak{V}_{\text{res}}$ irrelevant to the dynamics. Importantly, the existence of the commutator in Eq. (45) ensures nested commutators appearing in the SW transformation (44), such as $\mathfrak{S},\cdots,[\mathfrak{S},\mathfrak{L}_{0}]]$ with at least two $\mathfrak{S}$, or $\mathfrak{S},\cdots,[\mathfrak{S},\mathfrak{V}]]$ with at least one $\mathfrak{S}$, to be at least of order $\mathcal{O}(a^{-1})$. Hence, in the limit $a\to+\infty$, they vanish in the expansion of the SW transformation (44), leading to the simple generator
@@
[…]
@@
$\begin{array}[]{lll}&\displaystyle id\rho_{\mathbf{m},\mathbf{n}}/dt\\ &\displaystyle=[H_{S},\rho_{\mathbf{m},\mathbf{n}}]\\ &\displaystyle+\sum_{l}\Big((\nu_{l}-i\gamma_{l})m_{l}\rho_{\mathbf{m},\mathbf{n}}-(\nu_{l}+i\gamma_{l})n_{l}\rho_{\mathbf{m},\mathbf{n}}\Big)\\ &\displaystyle+\sum_{\mathbf{m}^{\prime}}\langle\mathbf{m}|X_{+}^{n}|\mathbf{m}^{\prime}\rangle s\rho_{\mathbf{m}^{\prime},\mathbf{n}}-\sum_{\mathbf{n}^{\prime}}\langle\mathbf{n}^{\prime}|X_{-}^{n}|\mathbf{n}\rangle\rho_{\mathbf{m},\mathbf{n}^{\prime}}s\\ &\displaystyle+\sum_{\mathbf{n}^{\prime}}\langle\mathbf{n}^{\prime}|X_{-,d^{\dagger}}^{n}|\mathbf{n}\rangle s\rho_{\mathbf{m},\mathbf{n}^{\prime}}-\sum_{\mathbf{m}^{\prime}}\langle\mathbf{m}|X_{+,d}^{n}|\mathbf{m}^{\prime}\rangle\rho_{\mathbf{m}^{\prime},\mathbf{n}}s\\ &\displaystyle+\sum_{r=1}^{n-1}\binom{n}{r}\sum_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}\bigg[\langle\mathbf{m}|X_{+}^{n-r}|\mathbf{m}^{\prime}\rangle\langle\mathbf{n}^{\prime}|X_{-,d^{\dagger}}^{n-r}|\mathbf{n}\rangle s\rho_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}\\ &\displaystyle-\langle\mathbf{m}|X_{+,d}^{n-r}|\mathbf{m}^{\prime}\rangle\langle\mathbf{n}^{\prime}|X_{-}^{n-r}|\mathbf{n}\rangle\rho_{\mathbf{m}^{\prime},\mathbf{n}^{\prime}}s\bigg],\end{array}$ (52)
@@
Appendix C Construction of the purified pseudomode model for simulating the fluorescence spectrum of quantum dots
Here we summarize the procedure for the construction of the purified pseudomode model for the quantum dot system described in section IV.2. The procedure consists of the following two steps.
  1. Step 1.
    Generate the spectral decomposition in Eq.(15) using the AAA algorithms. Other available algorithms can be found in Ref. [82]. This will produce the parameters $w_{l}$, $\nu_{l}$ and $\gamma_{l}$ required for the next step.
  2. Step 2.
    Substitute the obtain parameters $w_{l}$, $\nu_{l}$ and $\gamma_{l}$ into Eq.(23) and Eq. (24) to define the purified pseudomode model.
Once the model is build, one can simulate its dynamics to calculate physical quantities of interest under the initial condition $\rho_{\text{ppm}}(0)=\rho_{S}\otimes\rho_{\text{vac}}$ with $\rho_{\text{vac}}=\otimes_{\alpha}|0_{\alpha}\rangle\langle 0_{\alpha}|$ the vacuum state of all pseudomodes.